OpenStudy (dmezzullo):

Your lesson –– including a detailed explanation to the question provided along with at least 2 examples. Your reflection –– write a paragraph, of 3 or more sentences, as you reflect on this activity and your final product. You may use specific questions to guide your thoughts, if needed.

OpenStudy (dmezzullo):

@Kaitlyn_The_Math_wiz

OpenStudy (anonymous):

what activity? what?? lol what question? this might help

OpenStudy (dmezzullo):

this is 5.06 alg 1

OpenStudy (dmezzullo):

You must create a lesson that teaches another student how to solve the following inequality: –x + 3 > 7. Your main mission for this activity is to provide a detailed answer to the question below. Why must you flip the inequality symbol when you divide by a negative number? You know that if the symbol is not flipped you will end up with a false statement, but why is that so? Research and brainstorm (with your partner if you chose to do this collaboratively) to come up with this answer. Remember, you are teaching this to another student, so you must include detailed explanations, visuals, and at least two additional examples. You may submit your lesson in a word document, power point presentation, use audio/video tools, etc. If you are working with a partner, provide the student’s name, and outline what parts of the activity were completed by you, your partner, or where you worked together. Whether you are doing this activity independently or collaboratively you will submit the following to your instructor as assignment “05.06 Graphing Inequalities in Two Variables Activity." Your lesson –– including a detailed explanation to the question provided along with at least 2 examples. Your reflection –– write a paragraph, of 3 or more sentences, as you reflect on this activity and your final product. You may use specific questions to guide your thoughts, if needed.

OpenStudy (anonymous):

im not in algebra one, took it in public school. what exactly is it asking you to do?

OpenStudy (dmezzullo):

Up there explains it

OpenStudy (anonymous):

well first, do you know how to solve the inequality?

OpenStudy (dmezzullo):

Yes

OpenStudy (dmezzullo):

Here is the example i came upp with 1 sec

OpenStudy (dmezzullo):

-x-15=5 +15=+15 -x=20

OpenStudy (dmezzullo):

Would that be right?

OpenStudy (dmezzullo):

U sir seem like u can help me @ParthKohli

Parth (parthkohli):

Hello

OpenStudy (dmezzullo):

Can u help me

Parth (parthkohli):

The question is too long to read, and I am busy :-\

OpenStudy (dmezzullo):

ugh

OpenStudy (dmezzullo):

@hartnn